if x=-1 and y=2,what is the value of the expression 2x^3-3xy?
Answer by fractalier(2101)

To evaluate this, we merely plug in the values into
2x^3 - 3xy
and get
2(-1)^3 - 3(-1)(2) =
-2 + 6 =
4